# Spatially resolved broad line region in a quasar at z=4

### v1

@@ Acknowledgements.
We thank the staff of Paranal Observatory for their enthusiasm and help during the visitor runs in which the observations here were performed. This paper has benefitted from useful and lively discussion with many people on the BLR and outflows. In particular RD thanks B. Vollmer and K. Tristram, as well as the participants of the 2025 Vasto Accretion Meeting; RD and TS thank H. Netzer; JS thanks K. Inayoshi, J. Dai, and C. Jin for insightful suggestions on the BLR outflow modeling. JS is supported by The Fundamental Research Funds for the Central Universities, Peking University (7100604896) and the China Manned Space Program with grant no. CMS-CSST-2025-A09. LCH was supported by the National Science Foundation of China (12233001) and the China Manned Space Program (CMS-CSST-2025-A09). J.S-B. acknowledges the support received by the UNAM DGAPA-PAPIIT project AG-101025 and to the SECIHTI Ciencia de Frontera project CBF-2025-I-3033.References
- Abramowicz et al. (1988) Abramowicz, M. A., Czerny, B., Lasota, J. P., & Szuszkiewicz, E. 1988, ApJ, 332, 646
- Abuter et al. (2024) Abuter, R., Allouche, F., Amorim, A., et al. 2024, Nature, 627, 281
- Adamo et al. (2025) Adamo, A., Atek, H., Bagley, M. B., et al. 2025, Nature Astronomy [arXiv:2405.21054]
- Aird et al. (2015) Aird, J., Coil, A. L., Georgakakis, A., et al. 2015, MNRAS, 451, 1892
- Baskin & Laor (2018) Baskin, A. & Laor, A. 2018, MNRAS, 474, 1970
- Batiste et al. (2017) Batiste, M., Bentz, M. C., Raimundo, S. I., Vestergaard, M., & Onken, C. A. 2017, ApJ, 838, L10
- Bentz et al. (2013) Bentz, M. C., Denney, K. D., Grier, C. J., et al. 2013, ApJ, 767, 149
- Bentz et al. (2010) Bentz, M. C., Walsh, J. L., Barth, A. J., et al. 2010, ApJ, 716, 993
- Bertemes et al. (2025) Bertemes, C., Wylezalek, D., Rupke, D. S. N., et al. 2025, A&A, 693, A176
- Bischetti et al. (2017) Bischetti, M., Piconcelli, E., Vietri, G., et al. 2017, A&A, 598, A122
- Bogdán et al. (2024) Bogdán, Á., Goulding, A. D., Natarajan, P., et al. 2024, Nature Astronomy, 8, 126
- Bongiorno et al. (2014) Bongiorno, A., Maiolino, R., Brusa, M., et al. 2014, MNRAS, 443, 2077
- Boroson & Green (1992) Boroson, T. A. & Green, R. F. 1992, ApJS, 80, 109
- Chan & Krolik (2017) Chan, C.-H. & Krolik, J. H. 2017, ApJ, 843, 58
- Coatman et al. (2017) Coatman, L., Hewett, P. C., Banerji, M., et al. 2017, MNRAS, 465, 2120
- Collin et al. (2006) Collin, S., Kawaguchi, T., Peterson, B. M., & Vestergaard, M. 2006, A&A, 456, 75
- Czerny & Hryniewicz (2011) Czerny, B. & Hryniewicz, K. 2011, A&A, 525, L8
- Czerny et al. (2017) Czerny, B., Li, Y.-R., Hryniewicz, K., et al. 2017, ApJ, 846, 154
- Dalla Bontà et al. (2020) Dalla Bontà, E., Peterson, B. M., Bentz, M. C., et al. 2020, ApJ, 903, 112
- Davies et al. (2023) Davies, R., Absil, O., Agapito, G., et al. 2023, A&A, 674, A207
- Denney et al. (2009) Denney, K. D., Peterson, B. M., Pogge, R. W., et al. 2009, ApJ, 704, L80
- Du et al. (2016a) Du, P., Lu, K.-X., Hu, C., et al. 2016a, ApJ, 820, 27
- Du et al. (2016b) Du, P., Lu, K.-X., Zhang, Z.-X., et al. 2016b, ApJ, 825, 126
- Du & Wang (2019) Du, P. & Wang, J.-M. 2019, ApJ, 886, 42
- Du et al. (2018) Du, P., Zhang, Z.-X., Wang, K., et al. 2018, ApJ, 856, 6
- Elvis (2000) Elvis, M. 2000, ApJ, 545, 63
- Emmering et al. (1992) Emmering, R. T., Blandford, R. D., & Shlosman, I. 1992, ApJ, 385, 460
- Fabian (2012) Fabian, A. C. 2012, ARA&A, 50, 455
- Fan et al. (2023) Fan, X., Bañados, E., & Simcoe, R. A. 2023, ARA&A, 61, 373
- Faucher-Giguère & Quataert (2012) Faucher-Giguère, C.-A. & Quataert, E. 2012, MNRAS, 425, 605
- Ferrarese & Merritt (2000) Ferrarese, L. & Merritt, D. 2000, ApJ, 539, L9
- Förster Schreiber & Wuyts (2020) Förster Schreiber, N. M. & Wuyts, S. 2020, ARA&A, 58, 661
- Galbiati et al. (2023) Galbiati, M., Fumagalli, M., Fossati, M., et al. 2023, MNRAS, 524, 3474
- Gebhardt et al. (2000) Gebhardt, K., Bender, R., Bower, G., et al. 2000, ApJ, 539, L13
- Goad et al. (2012) Goad, M. R., Korista, K. T., & Ruff, A. J. 2012, MNRAS, 426, 3086
- GRAVITY+ Collaboration (2025) GRAVITY+ Collaboration. 2025, A&A, submitted
- GRAVITY Collaboration et al. (2017) GRAVITY Collaboration, Abuter, R., Accardo, M., et al. 2017, A&A, 602, A94
- GRAVITY+ Collaboration et al. (2022) GRAVITY+ Collaboration, Abuter, R., Allouche, F., et al. 2022, A&A, 665, A75
- GRAVITY Collaboration et al. (2021) GRAVITY Collaboration, Amorim, A., Bauböck, M., et al. 2021, A&A, 648, A117
- GRAVITY Collaboration et al. (2020) GRAVITY Collaboration, Amorim, A., Bauböck, M., et al. 2020, A&A, 643, A154
- GRAVITY Collaboration et al. (2024) GRAVITY Collaboration, Amorim, A., Bourdarot, G., et al. 2024, A&A, 684, A167
- Gravity Collaboration et al. (2024) Gravity Collaboration, Amorim, A., Bourdarot, G., et al. 2024, A&A, 690, A76
- GRAVITY Collaboration et al. (2018) GRAVITY Collaboration, Sturm, E., Dexter, J., et al. 2018, Nature, 563, 657
- Greene & Ho (2005) Greene, J. E. & Ho, L. C. 2005, ApJ, 630, 122
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
- Grier et al. (2013) Grier, C. J., Peterson, B. M., Horne, K., et al. 2013, ApJ, 764, 47
- Gültekin et al. (2009) Gültekin, K., Richstone, D. O., Gebhardt, K., et al. 2009, ApJ, 698, 198
- Heckman & Best (2014) Heckman, T. M. & Best, P. N. 2014, ARA&A, 52, 589
- Herrera-Camus et al. (2025) Herrera-Camus, R., González-López, J., Förster Schreiber, N., et al. 2025, A&A, 699, A80
- Hönig (2019) Hönig, S. F. 2019, ApJ, 884, 171
- Hopkins (2025) Hopkins, P. F. 2025, The Open Journal of Astrophysics, 8, 56
- Hopkins et al. (2025) Hopkins, P. F., Su, K.-Y., Murray, N., et al. 2025, The Open Journal of Astrophysics, 8, 48
- Horne et al. (2021) Horne, K., De Rosa, G., Peterson, B. M., et al. 2021, ApJ, 907, 76
- Hu et al. (2015) Hu, C., Du, P., Lu, K.-X., et al. 2015, ApJ, 804, 138
- Hu et al. (2008) Hu, C., Wang, J.-M., Ho, L. C., et al. 2008, ApJ, 683, L115
- Inayoshi et al. (2020) Inayoshi, K., Visbal, E., & Haiman, Z. 2020, ARA&A, 58, 27
- Jeon et al. (2024) Jeon, J., Bromm, V., Liu, B., & Finkelstein, S. 2024, in 45th COSPAR Scientific Assembly, Vol. 45, 1415
- Jiang & Dai (2024) Jiang, Y.-F. & Dai, L. 2024, arXiv e-prints, arXiv:2408.16856
- Jiang et al. (2019) Jiang, Y.-F., Stone, J. M., & Davis, S. W. 2019, ApJ, 880, 67
- Juodžbalis et al. (2025) Juodžbalis, I., Maiolino, R., Baker, W. M., et al. 2025, arXiv e-prints, arXiv:2504.03551
- Kaspi et al. (2021) Kaspi, S., Brandt, W. N., Maoz, D., et al. 2021, ApJ, 915, 129
- Kim et al. (2024) Kim, S. J., Goto, T., Ling, C.-T., et al. 2024, MNRAS, 527, 5525
- King & Pounds (2015) King, A. & Pounds, K. 2015, ARA&A, 53, 115
- King et al. (2011) King, A. R., Zubovas, K., & Power, C. 2011, MNRAS, 415, L6
- Kocevski et al. (2023) Kocevski, D. D., Onoue, M., Inayoshi, K., et al. 2023, ApJ, 954, L4
- Kokorev et al. (2023) Kokorev, V., Fujimoto, S., Labbe, I., et al. 2023, ApJ, 957, L7
- Kormendy & Ho (2013) Kormendy, J. & Ho, L. C. 2013, ARA&A, 51, 511
- Kovačević et al. (2010) Kovačević, J., Popović, L. Č., & Dimitrijević, M. S. 2010, ApJS, 189, 15
- Kudoh et al. (2023) Kudoh, Y., Wada, K., Kawakatu, N., & Nomura, M. 2023, ApJ, 950, 72
- Kudoh et al. (2024) Kudoh, Y., Wada, K., Kawakatu, N., & Nomura, M. 2024, ApJ, 977, 48
- Kuhn et al. (2024) Kuhn, L., Shangguan, J., Davies, R., et al. 2024, A&A, 684, A52
- Kwan & Krolik (1981) Kwan, J. & Krolik, J. H. 1981, ApJ, 250, 478
- La Mura et al. (2007) La Mura, G., Popović, L. Č., Ciroi, S., Rafanelli, P., & Ilić, D. 2007, ApJ, 671, 104
- Lai et al. (2023) Lai, S., Wolf, C., Onken, C. A., & Bian, F. 2023, MNRAS, 521, 3682
- Lambrides et al. (2024) Lambrides, E., Garofali, K., Larson, R., et al. 2024, arXiv e-prints, arXiv:2409.13047
- Larson et al. (2023) Larson, R. L., Finkelstein, S. L., Kocevski, D. D., et al. 2023, ApJ, 953, L29
- Le Fèvre et al. (2020) Le Fèvre, O., Béthermin, M., Faisst, A., et al. 2020, A&A, 643, A1
- Lira et al. (2018) Lira, P., Kaspi, S., Netzer, H., et al. 2018, ApJ, 865, 56
- Loeb & Barkana (2001) Loeb, A. & Barkana, R. 2001, ARA&A, 39, 19
- Long et al. (2023) Long, K., Dexter, J., Cao, Y., et al. 2023, ApJ, 953, 184
- Lupi et al. (2024) Lupi, A., Trinca, A., Volonteri, M., Dotti, M., & Mazzucchelli, C. 2024, A&A, 689, A128
- Maiolino et al. (2024) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2024, A&A, 691, A145
- Maoz et al. (1993) Maoz, D., Netzer, H., Peterson, B. M., et al. 1993, ApJ, 404, 576
- Marziani et al. (2017) Marziani, P., Negrete, C. A., Dultzin, D., et al. 2017, Frontiers in Astronomy and Space Sciences, 4, 16
- Matsuoka et al. (2007) Matsuoka, Y., Oyabu, S., Tsuzuki, Y., & Kawara, K. 2007, ApJ, 663, 781
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- McConnell & Ma (2013) McConnell, N. J. & Ma, C.-P. 2013, ApJ, 764, 184
- Mehta et al. (2024) Mehta, D., Regan, J. A., & Prole, L. 2024, The Open Journal of Astrophysics, 7, 107
- Mizumoto et al. (2021) Mizumoto, M., Nomura, M., Done, C., Ohsuga, K., & Odaka, H. 2021, MNRAS, 503, 1442
- Murray et al. (1995) Murray, N., Chiang, J., Grossman, S. A., & Voit, G. M. 1995, ApJ, 451, 498
- Nemmen & Brotherton (2010) Nemmen, R. S. & Brotherton, M. S. 2010, MNRAS, 408, 1598
- Netzer (2019) Netzer, H. 2019, MNRAS, 488, 5185
- Netzer (2020) Netzer, H. 2020, MNRAS, 494, 1611
- Netzer & Davidson (1979) Netzer, H. & Davidson, K. 1979, MNRAS, 187, 871
- Osterbrock (1978) Osterbrock, D. E. 1978, Proceedings of the National Academy of Science, 75, 540
- Osterbrock (1989) Osterbrock, D. E. 1989, Astrophysics of gaseous nebulae and active galactic nuclei (University Science Books)
- Pancoast et al. (2014) Pancoast, A., Brewer, B. J., & Treu, T. 2014, MNRAS, 445, 3055
- Park et al. (2022) Park, D., Barth, A. J., Ho, L. C., & Laor, A. 2022, ApJS, 258, 38
- Park et al. (2013) Park, D., Woo, J.-H., Denney, K. D., & Shin, J. 2013, ApJ, 770, 87
- Prince et al. (2023) Prince, R., Zajaček, M., Panda, S., et al. 2023, A&A, 678, A189
- Ramolla et al. (2018) Ramolla, M., Haas, M., Westhues, C., et al. 2018, A&A, 620, A137
- Ramos Almeida & Ricci (2017) Ramos Almeida, C. & Ricci, C. 2017, Nature Astronomy, 1, 679
- Reines et al. (2013) Reines, A. E., Greene, J. E., & Geha, M. 2013, ApJ, 775, 116
- Richards et al. (2006) Richards, G. T., Lacy, M., Storrie-Lombardi, L. J., et al. 2006, ApJS, 166, 470
- Robertson (2022) Robertson, B. E. 2022, ARA&A, 60, 121
- Runnoe et al. (2013) Runnoe, J. C., Brotherton, M. S., Shang, Z., & DiPompeo, M. A. 2013, MNRAS, 434, 848
- Saglia et al. (2016) Saglia, R. P., Opitsch, M., Erwin, P., et al. 2016, ApJ, 818, 47
- Santos et al. (2025) Santos, D. J. D., Shimizu, T., Davies, R., et al. 2025, A&A, 696, A30
- Scholtz et al. (2025) Scholtz, J., Maiolino, R., D’Eugenio, F., et al. 2025, A&A, 697, A175
- Shen et al. (2024) Shen, Y., Grier, C. J., Horne, K., et al. 2024, ApJS, 272, 26
- Shen & Ho (2014) Shen, Y. & Ho, L. C. 2014, Nature, 513, 210
- Shimizu et al. (2019) Shimizu, T. T., Davies, R. I., Lutz, D., et al. 2019, MNRAS, 490, 5860
[…]
- Somerville & Davé (2015) Somerville, R. S. & Davé, R. 2015, ARA&A, 53, 51
- Speagle (2020) Speagle, J. S. 2020, MNRAS, 493, 3132
- Suh et al. (2025) Suh, H., Scharwächter, J., Farina, E. P., et al. 2025, Nature Astronomy, 9, 271
- Sulentic & Marziani (2015) Sulentic, J. & Marziani, P. 2015, Frontiers in Astronomy and Space Sciences, 2, 6
- Sulentic et al. (2011) Sulentic, J., Marziani, P., & Zamfir, S. 2011, Baltic Astronomy, 20, 427
- Tacconi et al. (2020) Tacconi, L. J., Genzel, R., & Sternberg, A. 2020, ARA&A, 58, 157
- Temple et al. (2023) Temple, M. J., Matthews, J. H., Hewett, P. C., et al. 2023, MNRAS, 523, 646
- Tsuzuki et al. (2006) Tsuzuki, Y., Kawara, K., Yoshii, Y., et al. 2006, ApJ, 650, 57
- Übler et al. (2024) Übler, H., D’Eugenio, F., Perna, M., et al. 2024, MNRAS, 533, 4287
- Übler et al. (2023) Übler, H., Maiolino, R., Curtis-Lake, E., et al. 2023, A&A, 677, A145
- Vayner et al. (2024) Vayner, A., Zakamska, N. L., Ishikawa, Y., et al. 2024, ApJ, 960, 126
- Véron-Cetty et al. (2004) Véron-Cetty, M. P., Joly, M., & Véron, P. 2004, A&A, 417, 515
- Vestergaard & Peterson (2006) Vestergaard, M. & Peterson, B. M. 2006, ApJ, 641, 689
- Vietri et al. (2020) Vietri, G., Mainieri, V., Kakkad, D., et al. 2020, A&A, 644, A175
- Vietri et al. (2018) Vietri, G., Piconcelli, E., Bischetti, M., et al. 2018, A&A, 617, A81
- Vollmer et al. (2018) Vollmer, B., Schartmann, M., Burtscher, L., et al. 2018, A&A, 615, A164
- Ward et al. (2024) Ward, S. R., Costa, T., Harrison, C. M., & Mainieri, V. 2024, MNRAS, 533, 1733
- Watarai (2006) Watarai, K.-y. 2006, ApJ, 648, 523
- Wolf et al. (2024) Wolf, C., Lai, S., Onken, C. A., et al. 2024, Nature Astronomy, 8, 520
- Woo et al. (2024) Woo, J.-H., Wang, S., Rakshit, S., et al. 2024, ApJ, 962, 67
- Woo et al. (2015) Woo, J.-H., Yoon, Y., Park, S., Park, D., & Kim, S. C. 2015, ApJ, 801, 38
- Wright (2006) Wright, E. L. 2006, PASP, 118, 1711
- Wylezalek et al. (2022) Wylezalek, D., Vayner, A., Rupke, D. S. N., et al. 2022, ApJ, 940, L7
- Yong et al. (2020) Yong, S. Y., Webster, R. L., King, A. L., et al. 2020, MNRAS, 491, 1320
- Yong et al. (2017) Yong, S. Y., Webster, R. L., King, A. L., et al. 2017, PASA, 34, e042
- Zubovas & King (2012) Zubovas, K. & King, A. 2012, ApJ, 745, L34
- Zubovas & Nardini (2020) Zubovas, K. & Nardini, E. 2020, MNRAS, 498, 3633
@@ Appendix A Details of the model fit
[H] | flux scaling for peak of line profile | G(0.14,0.02) | 0.132 | -0.005,+0.005 |
The parameters for the best-fit model, and their 1- ranges, are given in Table 4. Here they have been grouped to make their role in the model easier to follow. The model has been fitted to the H and H lines simultaneously. As described in Kuhn et al. (2024), in practice this is done using two models in which all the geometry parameters as well as the global line emission parameters are tied to be the same in both. The only parameters that can differ between the models are those listed specifically for the H versus H line properties. These correspond to the shape parameter , the radial scaling parameter , and the peak flux scaling .
Histograms of the posterior distributions for the parameters are shown in Fig. 9, indicating that their ranges are generally well constrained by the data. As was done by Kuhn et al. (2024), we used the Python package dynesty (Speagle 2020) to fit the data with a nested sampling algorithm. Compared to the typical Markov Chain Monte Carlo algorithm, this method is powerful in terms of handling complex models with many (e.g. ) parameters and a potentially multimodal posterior distribution. We used the dynamic nested sampler (DynamicNestedSampler) which better estimates the likelihood function by re-sampling the posterior function a few more times after the “baseline run.” We used 1000 live points for the baseline run and added 500 points for each of the re-samplings until the default stop criteria was met. We adopted the random walk algorithm (rwalk) to sample the prior space and used the multi-ellipse method (multi) to create the nest boundaries. We adopted the default values for all the remaining options of dynesty. The reported fitting ends up with 2000 live points and effective samples, which is a statistically sufficient sampling of the posterior space. A corner plot of selected parameters is presented in Fig. 10. In addition to the high-probability peak, some low-probability solutions appear due to the very detailed sampling.

### v2

@@ Acknowledgements.
GRAVITY+ is developed in a collaboration by the Max Planck Institute for extraterrestrial Physics, the Institute National des Sciences de l’Univers du CNRS (INSU) with its institutes LIRA / Paris Observatory-PSL, IPAG / Grenoble Observatory, Lagrange / Côte d’Azur Observatory and CRAL / Lyon Observatory, the Max Planck Institute for Astronomy, the University of Cologne, the CENTRA - Centro de Astrofisica e Gravitação, the University of Southampton, the Katholieke Universiteit Leuven, University College Dublin, Universidad Nacional Autónoma de México and the European Southern Observatory.We thank the staff of Paranal Observatory for their enthusiasm and help during the visitor runs in which the observations here were performed. This paper has benefitted from useful and lively discussion with many people on the BLR and outflows. In particular RD thanks B. Vollmer and K. Tristram, as well as the participants of the 2025 Vasto Accretion Meeting; RD and TS thank H. Netzer; JS thanks K. Inayoshi, J. Dai, and C. Jin for insightful suggestions on the BLR outflow modeling. JS is supported by The Fundamental Research Funds for the Central Universities, Peking University (7100604896) and the China Manned Space Program with grant no. CMS-CSST-2025-A09. LCH was supported by the National Science Foundation of China (12233001) and the China Manned Space Program (CMS-CSST-2025-A09). J.S-B. acknowledges the support received by the UNAM DGAPA-PAPIIT project AG-101025 and to the SECIHTI Ciencia de Frontera project CBF-2025-I-3033. SH acknowledges support through UK Research and Innovation (UKRI) under the UK government’s Horizon Europe Funding Guarantee (EP/Z533920/1, selected in the 2023 ERC Advanced Grant round) and an STFC Small Award (ST/Y001656/1). This work has been supported by the French Agence Nationale de al Recherche (ANR) through contracts AGN_Melba ANR-21-CE31-001 and ExoVLTI ANR-21-CE31-0017. We also thank the Fundação para a Ciência e Tecnologia (FCT), Portugal, for the financial support to the Center for Astrophysics and Gravitation (CENTRA/IST/ULisboa) through grant No. UID/PRR/00099/2025 (https://doi.org/10.54499/UID/PRR/00099/2025) and grant No. UID/00099/2025 (https://doi.org/10.54499/UID/00099/2025). The authors thank D. McConnell for providing the RACS measurements. These data were obtained from Inyarrimanha Ilgari Bundara / the Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility.
@@ Appendix A Details of the model fit
The parameters for the best-fit model, and their 1- ranges, are given in Table 4. Here they have been grouped to make their role in the model easier to follow. The model has been fitted to the H and H lines simultaneously. As described in kuh24, in practice this is done using two models in which all the geometry parameters as well as the global line emission parameters are tied to be the same in both. The only parameters that can differ between the models are those listed specifically for the H versus H line properties. These correspond to the shape parameter , the radial scaling parameter , and the peak flux scaling .
Histograms of the posterior distributions for the parameters are shown in Fig. 8, indicating that their ranges are generally well constrained by the data. As was done by kuh24, we used the Python package dynesty (spe2020) to fit the data with a nested sampling algorithm. Compared to the typical Markov Chain Monte Carlo algorithm, this method is powerful in terms of handling complex models with many (e.g. ) parameters and a potentially multimodal posterior distribution. We used the dynamic nested sampler (DynamicNestedSampler) which better estimates the likelihood function by re-sampling the posterior function a few more times after the “baseline run.” We used 1000 live points for the baseline run and added 500 points for each of the re-samplings until the default stop criteria was met. We adopted the random walk algorithm (rwalk) to sample the prior space and used the multi-ellipse method (multi) to create the nest boundaries. We adopted the default values for all the remaining options of dynesty. The reported fitting ends up with 2000 live points and effective samples, which is a statistically sufficient sampling of the posterior space. A corner plot of selected parameters is presented in Fig. 9. In addition to the high-probability peak, some low-probability solutions appear due to the very detailed sampling.
@@
| [H] | flux scaling for peak of line profile | G(0.14,0.02) | 0.132 | -0.005,+0.005 | |